OpenStudy (anonymous):

URGENT:Assume a 150 pound person burns 120 calories after 2 hours of sleep and 240 calories after 4 hours of sleep. Part 1 (2 points): Use calories as the y-coordinate and hours as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 8 hours of sleep? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

had u found the slope?

OpenStudy (anonymous):

1/60?

OpenStudy (anonymous):

yes so in 0 hours he burns 0 calories u can use the formula for the line y-y0 =m(x-x0) with (yo,xo) = (0,0)

OpenStudy (anonymous):

sorry the slope is 60

OpenStudy (anonymous):

wait what do i do...?

OpenStudy (anonymous):

One equation for a straight line is: \[y-y_{1}=m(x-x_{1})\]where: m=slope x_1 is the x coordinate of a point on the line and y_1 is the y coordinate of a point on the line. You know the points (2,120) and (4,240), so you can plug any of them in. The slope is going to be the change in y of these two points over the change in x, or (240-120)/(4-2)=60. The slope represents the calories burned per hour. Plug these in and you get \[y-120=60(x-2)\] Simplify and you'll find \[y=60x\] To find the calories burned in 8 hours, plug in 8 for x, and you get 480.

OpenStudy (anonymous):

so thats part 1?

OpenStudy (anonymous):

To recap: For Part 1: The slope represents the change in y (calories) per change in x (hours) so the slope is the amount of calories burned per hour. Part 2: y=60x Part 3: 480 calories See my previous answer which has full explanations and work shown.

OpenStudy (anonymous):

OH thanks so much!!

OpenStudy (anonymous):

No problem, glad to help.